OpenStudy (herro):

Matt looks at the architectural plan of a four-walled room in which the walls meet each other at right angles. The length of one wall in the plan is 19 inches. The length of the diagonal of the floor of the room in the plan is approximately 20.62 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work. is this the right answer??No, the floor plan is not a square it is a rectangle because the length is 19 inches and the width is 20.62 inches

OpenStudy (herro):

@TheSmartOne

OpenStudy (herro):

@Nnesha

OpenStudy (asc.bchs):

You are right that it is not a square, but your reason was wrong. It says the diagnosis is 20.62 inches, not the width. You can check if it is a square with the Pythagorean theorem. A^2 + B^2 = C^2 Assume it is a square. A & B would both be 19 because one side was 19 inches as stated. Plug it in 19^2 + 19^2 = C^2 361 + 361 = C^2 722 = C^2 Square root both sides 26.87 = C If the room were a square, the diagonal would be 26.87 inches. You can also work in reverse, using 20.62 as C. 19^2 + B^2 = 20.62^2 361 + B^2 = 425.1844 B^2 = 64.1844 B = 8.0115... The sides are 19 inches and about 8 inches, therefore the room is not a square.

OpenStudy (herro):

thank you so much @asc.bchs

OpenStudy (herro):

can u help with a few more??

OpenStudy (asc.bchs):

Sure